### v1

# MaskedManipulator: Versatile Whole-Body Control for Loco-Manipulation

Abstract.
Humans interact with their world while leveraging precise full-body control to achieve versatile goals. This versatility allows them to solve long-horizon, underspecified problems, such as placing a cup in a sink, by seamlessly sequencing actions like approaching the cup, grasping, transporting it, and finally placing it in the sink. Such goal-driven control can enable new procedural tools for animation systems, enabling users to define partial objectives while the system naturally “fills in” the intermediate motions. However, while current methods for whole-body dexterous manipulation in physics-based animation achieve success in specific interaction tasks, they typically employ control paradigms (e.g., detailed kinematic motion tracking, continuous object trajectory following, or direct VR teleoperation) that offer limited versatility for high-level goal specification across the entire coupled human-object system. To bridge this gap, we present MaskedManipulator, a unified and generative policy developed through a two-stage learning approach. First, our system trains a tracking controller to physically reconstruct complex human-object interactions from large-scale human mocap datasets. This tracking controller is then distilled into MaskedManipulator, which provides users with intuitive control over both the character’s body and the manipulated object. As a result, MaskedManipulator enables users to specify complex loco-manipulation tasks through intuitive high-level objectives (e.g., target object poses, key character stances), and MaskedManipulator then synthesizes the necessary full-body actions for a physically simulated humanoid to achieve these goals, paving the way for more interactive and life-like virtual characters.
copyright: none
Figure 1. MaskedManipulator enables physics-based humanoids to perform intricate, object interactions from sparse spatio-temporal goals.
1. Introduction
Generating versatile humanoid agents capable of autonomous and controllable behaviors is a shared challenge for both character animation and robotics. While recent approaches have significantly advanced locomotion capabilities for simulated (Tessler et al., 2024a; Wu et al., 2025) and real-world humanoids (Ji et al., 2024; Allshire et al., 2025; Ze et al., 2025), synthesizing controllers that can skillfully interact with the environment – especially with small, dynamic objects requiring precise dexterous manipulation – remains a daunting hurdle. A key challenge lies in how to effectively translate high-level human intent into low-level motor commands. For example, walking to an object, picking it up, inspecting the object, and then placing it back down on a table.
We aim to build a versatile controller that seamlessly combines full-body locomotion with dexterous manipulation, which can also interpret such diverse command inputs. However, there exits an inherent tension: the need for broad task flexibility versus the need for exacting physical precision. On one hand, achieving versatility requires controllers capable of understanding abstract user intents, like sparsely defined target coordinates (e.g., “reach this point”) or high-level object goals (e.g., “place cup here”). These often underspecified goals present vast solution spaces, posing substantial challenges for control and learning techniques like reinforcement learning (RL). On the other hand, successful manipulation, particularly with common objects, demands meticulous precision. Even small deviations during physical interaction – grasping, repositioning, or placing – can lead to instability, object dropping, or complete task failure, often without recovery (Luo et al., 2024b). While prior work has tackled aspects of this challenge, existing methods typically excel at either sophisticated manipulation (Lin et al., 2024; Iyer et al., 2024) or versatile body control (e.g., via goal-conditioning (Tessler et al., 2024a)). Achieving unified versatile control that encompasses both body and object goals, and gracefully balances this flexibility-precision trade-off, remains an open problem.
We propose a unified control framework for diverse and precise loco-manipulation tasks, responding to inputs ranging from detailed kinematic targets (e.g., tracking the hand positions via teleoperation) to sparse, high-level goals (like a desired position of an object). Our framework extends spatio-temporal goal-conditioning to encompass both the humanoid’s body parts and the manipulated object. To master the precise execution required by human-object interactions, our first stage (MimicManipulator) leverages human motion capture (e.g., GRAB dataset (Taheri et al., 2020)). The motion capture data demonstrates successful interaction strategies, such as sequences of grasping, placing, and regrasping objects, and hand-to-hand transfers. This physics-based tracker (MimicManipulator), trained with RL in a fully observable setting, learns the necessary actions to accurately reconstruct these intricate behaviors. Then, to enable versatile control from sparse goals, MimicManipulator’s expert interaction knowledge is distilled into our MaskedManipulator policy. By learning from MimicManipulator, MaskedManipulator becomes capable of producing human-like and diverse behaviors in response to under-specified goals for both body parts and objects (e.g., “bring the object to coordinate $(x,y,z)$”), effectively addressing the multi-solution challenge of versatile control.
The central contributions of this work are:
  1. (1)
    A novel framework for versatile loco-manipulation that leverages human demonstrations to produce diverse, physically plausible, and human-like solutions to underspecified tasks involving coupled human-object interactions.
  2. (2)
    MimicManipulator: A physics-based motion tracker that accurately infers actions and reconstructs dexterous human loco-manipulation sequences from reference motion data.
  3. (3)
    MaskedManipulator: A unified generative loco-manipulation policy that enables diverse spatio-temporal goal-conditioning for both humanoid body parts and manipulated objects.
2. Related Work
Our work builds upon advances in demonstration-driven control, object manipulation, and versatile goal-conditioned policies for physics-based characters.
Demonstration-driven control.
Traditional reinforcement learning (RL) for character control often requires meticulous reward engineering. Imitation learning (IL) offers a powerful alternative by leveraging expert demonstrations. DeepMimic (Peng et al., 2018) showed RL can robustly replicate reference motions, a paradigm extended to object manipulation by works like (Yu et al., 2025) for dexterous sequences and InterMimic (Xu et al., 2025) for retargeting interactions across morphologies. Our MimicManipulator stage is inspired by these approaches, using RL to train a single policy to precisely track the entire diverse human loco-manipulation demonstrations from GRAB (Taheri et al., 2020), thereby recovering actions and ensuring physical plausibility.
Object manipulation and grasping
Integrating object manipulation significantly increases control complexity. Some approaches achieve manipulation without reference data via hand-crafted rewards and large-scale RL (Akkaya et al., 2019; Lin et al., 2024). More recently, learning-based methods using demonstrations have enabled full-body motion for object-trajectory tracking (Luo et al., 2024a). While these methods show progress in specific manipulation aspects, our work focuses on versatile, goal-conditioned manipulation from human data. MaskedManipulator uniquely allows explicit conditioning on the future state of the object itself as a goal, alongside humanoid body part goals. This enables synthesizing novel, full-body loco-manipulation behaviors from sparse constraints, offering a more flexible control paradigm.
Versatile and goal-conditioned control
A trend in robotics and animation involves hierarchical control, where high-level systems specify abstract goals (e.g., walk targets (Peng et al., 2022), grasp poses (Li et al., 2024), 2D paths (Li et al., 2025)) for low-level execution, akin to System 1/System 2 reasoning (Kahneman, 2011). MaskedMimic (Tessler et al., 2024a) significantly advanced System1 versatility by framing character control as motion inpainting, allowing a unified model to synthesize diverse motions from sparse future joint targets. Our work extends this by enabling explicit spatio-temporal goal-conditioning on both humanoid body parts and the manipulated object within MaskedManipulator. This provides direct, versatile control over the entire coupled human-object system, crucial for sophisticated loco-manipulation tasks.
3. Preliminaries
We formulate our problem within the framework of goal-conditioned Markov Decision Process (MDP) (Puterman, 2014), defined by a tuple $M=(S,G,A,P,R,\gamma)$, consisting of states $S$, goals $G$, actions $A$, transition dynamics $P$, a reward function $R$, and a discount factor $\gamma$. The objective is to learn a policy $\pi(a_{t}|s_{t},g_{t})$ that selects an action $a_{t}$ based on the current state $s_{t}$ and a task-specific goal $g_{t}$.
The state of the humanoid’s $J$ links at time $t$ is described by their individual 3D positions $p_{j,t}\in\mathbb{R}^{3}$ orientations $\theta_{j,t}$, linear $v_{j,t}$ and angular $\omega_{j,t}$ velocities. Each orientation $\theta_{j,t}$ is represented using a continuous 6D rotation vector (Zhou et al., 2019). All link positions and orientations constitute the full body kinematic state, denoted $q_{t}=(\{p_{j,t}\}_{j=1}^{J},\{\theta_{j,t}\}_{j=1}^{J})$. The corresponding linear and angular velocities collectively form $\dot{q}_{t}$. When interacting with the scene, a body part may be subjected to contact forces, which we denote by $c_{j,t}$. Object states $q^{\text{obj}}_{t}$ are similarly defined by 3D position $p^{\text{obj}}_{t}$, 6D orientation $\theta^{\text{obj}}_{t}$, velocities $\dot{q}^{\text{obj}}_{t}$, and contact forces $c^{\text{obj}}_{t}$.
Reference kinematic and contact quantities, sourced from Motion Capture (MoCap) data or a target trajectory generator, are denoted with a hat accent (e.g., $\hat{q}_{t},\hat{\dot{q}}_{t},\hat{q}^{\text{obj}}_{t},\hat{c}_{t},\hat{c}^{\text{obj}}_{t}$). In contrast to the simulated information, the reference $\hat{c}_{t}$ are contact indicators. Quantities without this accent refer to states within our physics simulation.
4. Method
Our approach for generating versatile humanoid loco-manipulation capabilities is structured around a two-stage learning process. Human motion capture data provides rich kinematic descriptions of complex interactions but typically lacks the underlying low-level motor commands (actions) required to drive a physics-based character. Our first stage, MimicManipulator, addresses these challenges by training a high-fidelity motion tracker within a physics-based simulation. Operating in a full-information setting with respect to the reference kinematics, MimicManipulator learns not only to recover the necessary actions but also to physically reconstruct the target loco-manipulation sequences, thereby ensuring dynamic feasibility and emphasizing the nuances of object handling. Subsequently, the robust interaction skills learned by MimicManipulator is distilled into a versatile control policy, MaskedManipulator. This second stage yields a policy that can generate novel, physically plausible behaviors by dynamically conditioning on spatio-temporal goals specified for various parts of the humanoid body or the manipulated object itself. The following sections detail the objectives, architecture, and training procedures for each of the two stages.
4.1. MimicManipulator
The first stage focuses on reproducing the intricate behaviors required for dexterous loco-manipulation. We train a single policy $\pi_{\text{track}}$, a motion-tracker denoted by MimicManipulator, to accurately reproduce the complex loco-manipulation sequences from the processed GRAB dataset. The reference kinematic data provides a rich description of the desired high-level behavior but does not contain the low-level motor commands (actions $a_{t}$) necessary to execute these motions in a physically simulated environment. Therefore, we formulate this task as a reinforcement learning problem where $\pi_{\text{track}}$ learns to infer these actions by imitating the reference motion.
4.1.1. Task design
(a) Pre-contact
(b) During contact
(c) Post-contact
Figure 2. Phased contact reward for precise manipulation. We design a three-stage contact reward: (a) Approach: Tracks the reference motion while aligning the hand’s path relative to the object surface. (b) Engagement: Ensures critical contacts are maintained according to the reference. (c) Release: Promotes a smooth and timely object disengagement mirroring the demonstration. The spheres illustrate the reference joints and the arrows illustrate vectors to the closest point on the object’s surface.
Observation.
To enable object interactions, the observation $s_{t}$ comprises of the character’s proprioception $q_{t}$, object information $q^{\text{obj}}_{t}$ and character-object relational features. All components are expressed in a character’s local coordinate frame and aligned with it’s current root heading (yaw) for policy invariance. The character-object relational features capture the spatial relationship between the humanoid and the object, by providing the vector (direction and distance) from each body part to the object surface. Object structure: Precise interaction requires awareness of the object geometry. To accomodate this, we also provide a Basis-Point-Set (Prokudin et al., 2019, BPS) representation, which represents all objects by the surface distances to a pre-sampled set of points. Compared to point-cloud representations, the pre-sampled and fixed BPS allow processing using simpler architectures (e.g., MLP), allowing faster training and inference, while still being expressive. Future reference trajectory (goal): To reproduce the target motions, MimicManipulator observes the the goal future poses in the next frame $K=1$ and $K=30$ frames (1 second), denoted by $g_{t}^{\text{track}}$. Each target future pose observation $\{\hat{q}_{t+k}\Theta q_{t}\}=\{\hat{p}_{t+k}-p_{t},\hat{q}_{t+k}\Theta q_{t}\}$ is a future pose $\hat{q}_{t+k}$ normalized with respect to the current simulated pose $q_{t}$, where $\Theta$ represents the inverse quaternion. In addition, we provide $\hat{c}_{t+k}$ to indicate which body parts should have contact with the object.
Reward and termination.
Successfully reconstructing long and intricate human-object interaction sequences requires high precision. Minor inaccuracies during critical interaction phases (e.g., grasping, placing) can compound, leading to failures much later in the motion. This delayed consequence creates a challenging credit assignment problem for reinforcement learning.
Reference demonstrations implicitly encode successful, precise interaction strategies. For example, where to grab the object in order to naturally reach a target orientation, or to successfully perform a hand-to-hand transfer. We leverage this data and design reward $R_{\text{track}}$ and termination conditions to synergistically guide the MimicManipulator agent towards accurate and physically plausible reconstructions.
The reward $R_{\text{track}}$ is a product of several components, ensuring high reward only when all aspects of the reference are accurately matched (Park et al., 2019; Won et al., 2020; Xu et al., 2025):
(1) $R_{\text{track}}=r^{\text{pose}}\cdot r^{\text{contact}}\cdot r^{\text{energy}}\cdot r^{\text{interaction}}\,.$
Each component $r^{(\cdot)}$ is an exponential of a negatively weighted error or cost (e.g., $r=\exp(-w\cdot\text{cost})$), valued between (0, 1]. $r^{\text{humanoid}}=r^{\text{ht}}\cdot r^{\text{hr}}$ and $r^{\text{obj}}=r^{\text{ot}}\cdot r^{\text{or}}$ components consider the translation and rotation errors for the human and object respectively. The energy component attempts to mitigate artifacts such as jitters. The interaction component aims to prevent human-object penetrations by encouraging a gentle approach and interaction.
The contact component, Figure 2, encourages the policy to replicate the reference contact patterns and timing throughout the interaction sequence. It guides the agent through: Approach & engagement: Smoothly guiding the end-effectors (e.g., hands) towards the target contact points on the object, matching the reference approach. Maintaining contact: Ensuring stable contact once established, consistent with the demonstration. Disengagement: Facilitating a clean and timely release, mirroring the reference disengagement.
We found that the multiplicative design helps learning precise, long-horizon loco-manipulation by strictly requiring competence in all elements of the complex motion. Complementing the reward design, we leverage early termination to define a feasibility envelope for the training phase (Peng et al., 2018; Luo et al., 2023; Tessler et al., 2024a). An episode is terminated if any body part deviates by ¿25cm from reference, or the object deviates ¿10cm from reference, or when there’s a prolonged unintended loss of contact for ¿10 consecutive frames during a contact phase, or if the character maintains contact ¿0.4sec after the reference has released.
Prioritized training for complex interactions.
We train a single policy to reconstruct the entire dataset. The reference sequences span a wide range of difficulty. This variation arises from both the object’s characteristics (e.g., grasping a mug by its handle is more demanding than grasping a banana) and the intricacy of the interaction itself (e.g., bimanual operations or tasks requiring precise tool-like dexterity). During training we periodically evaluate the agent’s performance and over-sample sequences in which it has failed (Luo et al., 2024a; Tessler et al., 2024a; Luo et al., 2023).
Figure 3. Object Retargeting for Morphological Differences. Transferring motion between characters of varying shapes can misalign human-object interactions (left). Our method leverages original contact data to retarget the object’s trajectory, preserving interaction consistency (right).
4.1.2. Data processing
(a) Deterministic
(b) C-VAE: A prior is trained to represent the set of plausible solutions for each state $s_{t}$ and sparse objectives $g_{t}^{\text{versatile}}$.
(c) Diffusion: Conditioned on the noisy action $a_{t}^{j}$ and the current diffusion timestep $j$.
Figure 4. MaskedManipulator architectures: We illustrate the inference procedure of the considered architectures.
The GRAB dataset consists of motion sequences collected across 10 different human subjects, differing by gender and size. This poses a challenge for training the controller, as it requires awareness and generalization across different morphologies. To simplfy the control problem, we use a single, canonical humanoid body shape (mean SMPL-X). The core difficulty lies in accurately retargeting the diverse source motions from different subjects to our canonical character while preserving the semantically salient characteristics of the original human-object interaction.
We propose a simple yet effective approach. First, we map the motion from source to target skeletons by transferring the local DoF rotations. Then, we retarget the object. Using original contact data from the aligned source motion, we optimize the object’s global position to best preserve these contact relationships with the canonical character’s end-effectors (additional details in the supplementary material and illustrated in Figure 3).
Despite retargeting, some motions remain unsuitable for physical reconstruction. We filter these, primarily due to: Complex bimanual interactions where our object-centric retargeting may not fully resolve hand-object-hand alignment, and interactions reliant on features our simplified humanoid model lacks (e.g., placing sunglasses on a detailed face). This data processing, including retargeting and filtering, yields a training set of 1007 motion sequences and a test set of 141 sequences (GRAB subject 10).
4.2. MaskedManipulator
The second stage of our framework distills the rich, physically-grounded interaction expertise learned by $\pi_{\text{track}}$ (from MimicManipulator) into a versatile, generative control policy, $\pi_{\text{versatile}}$ (our MaskedManipulator). The ultimate goal for MaskedManipulator is to produce diverse, human-like, and physically plausible full-body manipulation behaviors in response to sparse future goals. These goals might specify the desired pose for a hand, the target location for an object, or a full-body stance at a future time, effectively leaving many other aspects of the motion for the policy to “inpaint”.
Directly training such a versatile policy from scratch is exceptionally challenging. Sparse goals often lead to a misspecified objective with an overwhelmingly large space of potential solutions, posing significant exploration and credit assignment difficulties. However, MimicManipulator ($\pi_{\text{track}}$), having learned to reconstruct a wide array of complex human demonstrations, implicitly captures a distribution of high-quality, physically feasible solutions to various interaction sub-tasks. Our approach extends the methodology of MaskedMimic (Tessler et al., 2024a), which demonstrated that framing control as a “physics-based inpainting” problem – where the policy fills in missing parts of a motion sequence – can effectively learn versatile behaviors from diverse data.
Goal Specification and Distillation:
The core idea is to train MaskedManipulator ($\pi_{\text{versatile}}$) to predict the detailed actions that MimicManipulator ($\pi_{\text{track}}$) would execute to complete a motion, but conditioned only on the current state $s_{t}$ and a sparse set of future goals $g_{t}^{\text{versatile}}$. We employ an online teacher-student distillation process based on DAgger (Ross et al., 2011).
At each step, the student policy $\pi_{\text{versatile}}$ receives the current state $s_{t}$ and a randomly masked version of the future reference trajectory, $g_{t}^{\text{versatile}}$. This sparse goal $g_{t}^{\text{versatile}}$ specifies target future states (e.g., pose of a hand, object location, or full body stance) for a subset of entities at one or more future timesteps. All other future state information is masked. The student then predicts an action $a_{t}^{\text{versatile}}$ which is used to control the character in the simulation.
Concurrently, the teacher policy $\pi_{\text{track}}$ observes the same current state $s_{t}$ but with the unmasked, dense future reference trajectory $g_{t}^{\text{track}}$ (as used during its own training). The distillation objective is to minimize the L2 distance between the actions predicted by the student and the teacher:
(2) $\mathcal{L}_{\text{distill}}=-\log\pi_{\text{versatile}}\left(a_{t}^{\text{track}}|s_{t},g_{t}^{\text{versatile}}\right)$
We adapt the structured masking scheme from MaskedMimic (Tessler et al., 2024a) to define $g_{t}^{\text{versatile}}$, crucially extending it to include conditioning on the future pose of the manipulated object alongside humanoid body parts.
Policy Architecture and Training:
To accommodate a variable number of sparsely specified future goal constraints, each defined by a target pose and a future timestep, MaskedManipulator utilizes a transformer architecture (Vaswani et al., 2017). The input to the transformer is a sequence of tokens. Fixed-information inputs, such as current proprioception (joint states, root state), hand-to-object vectors, object and table pose, and object and table BPS, are jointly mapped to a shared token. Each specified future goal (e.g., target pose of hand $j$ at time $t+k$) is also encoded into its own token. Which each target future pose is represented using a unique token, they are generated using a shared encoder.
Given that a sparse goal $g_{t}^{\text{versatile}}$ can often be achieved through multiple valid and human-like motion sequences (i.e., the problem is multi-modal), we experiment with three classes of policy architectures for $\pi_{\text{versatile}}$, illustrated in Figure 4:
(1) Deterministic: This standard approach learns to predict a single, mean action based on the teacher’s demonstrations. While computationally efficient, this can average over the diverse solutions present in the data, potentially leading to less expressive or less performant behaviors when multiple valid options exist.
(2) Conditional Variational Autoencoder (C-VAE): To explicitly model the multi-modality of solutions, we employ a C-VAE architecture with a learned prior, drawing inspiration from (Rempe et al., 2021; Tessler et al., 2024a). The prior network takes the sparse goal $g_{t}^{\text{versatile}}$ and current state $s_{t}$ to predict a latent distribution $\mathcal{N}(\mu_{prior},\sigma_{prior})$ representing plausible general solutions. During training, a residual encoder provides the offset in the latent space, predicting the precise solution from the reference data. We extend the architecture in MaskedMimic. There, the encoder observes both the reference motion $g_{t}^{\text{track}}$ and the sparse goals $g_{t}^{\text{versatile}}$. As a result, the encoder is required to implicitly predict both the prior distribution and the offset. We simplify this design by providing the encoder with the prior’s output $\mathcal{N}(\mu_{prior},\sigma_{prior})$. We found it reduces parameter count and improves performance. During inference, sampling from the learned prior allows MaskedManipulator to generate diverse actions satisfying the sparse goal.
(3) Diffusion Policy: As an additional generative approach, we investigate a diffusion-based policy (Chi et al., 2023; Ho et al., 2020; Lu et al., 2025). Unlike C-VAEs, diffusion models do not require an explicit prior network or separate encoder for the posterior during training for their generative process. The policy is trained to denoise an action that is initially pure Gaussian noise, iteratively refining it over $N$ steps to produce a clean action $a_{t}^{N}$. This iterative refinement process is conditioned on the current state $s_{t}$ and sparse goal $g_{t}^{\text{versatile}}$. To provide the necessary conditioning for each denoising step $j$, we augment our transformer architecture with two additional input heads: one tokenizing the noisy action $a_{t}^{j}$ from the current step, and another representing the current denoising timestep $j$. At each step $j$, the policy predicts the less noisy action $a_{t}^{j-1}$ (or equivalently, the noise to remove). The final denoised action $a_{t}^{0}$ (conventionally, or $a_{t}^{N}$ if $N$ is the start of denoising from pure noise) is used to control the character. In addition to Diffusion using DAgger, we also experiment with a fully-offline version, where MaskedManipulator is trained using supervised learning on trajectories collected by MimicManipulator.
Both the C-VAE and Diffusion approaches are explicitly chosen for their capacity to model the inherent diversity and multi-modality in the solution space when presented with sparse goals. This empowers MaskedManipulator not merely to fulfill specified constraints, but to do so with a rich variety of human-like behaviors, reflecting the breadth of solutions learned by MimicManipulator.
Table 1. MimicManipulator: We ablate various design decisions, stripping another component and measuring the cumulative importance. E.g., the “Contact guidance” is without “Tight termination” and “Prioritized scenes”. We also compare with our implementation of InterMimic (Xu et al., 2025) (see Section 6.1).
Train Test
Sequence Success First Sequence Success MPJPE [mm] Tracking Length [s] Sequence Success First Sequence Success MPJPE [mm] Tracking Length [s]
[…]
(-) Tight termination 74.6% 91.2% 15.3 9.3 51.8% 80.1% 22.8 7
(-) Prioritized scenes 71.2% 89.3% 15.5 9.2 56% 73.8% 18.4 7
   (-) Contact guidance 69.4% 86.6% 16.4 9.2 47.5% 78% 25.5 6.5
InterMimic* (Xu et al., 2025) 11% 31.1% 42.2 6.2 8.5% 29.8% 50.5 4.9
5. Experiments
Experimental setup:
All experiments are conducted using the ProtoMotions framework (Tessler et al., 2024b) with IsaacLab (Mittal et al., 2023) as the underlying physics simulator. The simulation runs at 120 FPS, and with 4 physics decimation steps, our control policies operate at an effective rate of 30 FPS. To manage interpenetrations robustly, we set the simulation’s depenetration velocity to 100. To better approximate the contact dynamics of human soft tissue interacting with objects using rigid body simulation, we use a global friction coefficient of 1.5.
Training details:
All MimicManipulator tracking policies ($\pi_{\text{track}}$) are trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The MaskedManipulator variants ($\pi_{\text{versatile}}$) are trained using DAgger (Ross et al., 2011) with $\pi_{\text{track}}$ as the expert teacher. Each policy variant is trained for approximately 40,000 episodes, distributed across 8 NVIDIA A100 GPUs. Specific hyperparameters for PPO and the distillation process are detailed in the Appendix.
Evaluation metrics:
The loco-manipulation sequences from GRAB are often long and complex, involving multiple pick-and-place iterations, and hand-to-hand transfers. To comprehensively evaluate performance, we employ several metrics:
Full-Sequence Success Rate: The percentage of test sequences successfully completed from start to finish.
First-Interaction Success Rate: To isolate early interaction performance from later complexities, this measures success from the motion’s start until the second required contact with the object.
Mean Per Joint Position Error (MPJPE): The average Euclidean distance between the simulated joint positions and the reference motion’s joint positions, providing a measure of tracking fidelity.
Average Sequence Length: The average episode length until termination is encountered (or episode ends).
We define failure (leading to episode termination for success rate calculation) if any humanoid joint deviates by more than 50cm from its reference position, or if the manipulated object deviates by more than 25cm from its reference position (Luo et al., 2024a).
6. Results
6.1. MimicManipulator
We evaluate ManipulationMimic’s performance through ablation studies and quantitative comparisons, and provide qualitative examples of its capabilities.
Ablation study and quantitative analysis
Table 1 presents MimicManipulator’s performance on the GRAB dataset and the impact of key design choices. We ablated: (1) tight early termination criteria, (2) prioritized training on harder motions, and (3) the phased contact guidance reward. The results demonstrate that each component positively contributes to the overall tracking success, with the strict early termination yielding the most significant improvement by maintaining a tight feasibility envelope.
We also compare MimicManipulator with our implementation of InterMimic (Xu et al., 2025). For a fair comparison, we trained a single InterMimic model on the entire GRAB training set, similar to MimicManipulator, omitting InterMimic’s subject-specific policy distillation stages. In addition, we did not include physical state initialization as it introduced instabilities in dexterous manipulation tasks. MimicManipulator outperforms this adapted InterMimic baseline. We attribute InterMimic’s lower performance in this setting to its potentially looser tracking requirements and a lack of explicit object structure awareness (e.g., a BPS representation for contact guidance) which is crucial for precise interactions with diverse objects in GRAB.
Qualitative analysis
As illustrated in Figure 5 and the supplementary video, MimicManipulator successfully reconstructs a diverse range of complex, full-body interactions while preserving natural, human-like motion qualities. For instance, the policy can accurately simulate grasping a teapot by its handle to pour liquid, or picking up a hammer and striking a nail. These examples highlight its ability to learn nuanced, tool-specific behaviors. While MimicManipulator reliably reconstructs the majority of motions, occasional uncanny behaviors can be observed, particularly during the precise moments of contact initiation or release. These may stem from the discrete nature of the termination conditions related to contact, which can sometimes force abrupt transitions rather than perfectly smooth engagement or disengagement.
6.2. MaskedManipulator
We quantitatively evaluate MaskedManipulator ($\pi_{\text{versatile}}$), assessing its versatility, generalization, and long-horizon reasoning.
Quantitative analysis.
Teleoperation-Style Pose Matching: Table 2 compares the proposed MaskedManipulator architectures (Deterministic, C-VAE, Diffusion) on a teleoperation-like task. Here, the policy is conditioned on near-future goals for head and wrist poses, plus the object’s pose, aiming to accurately achieve these combined configurations. While training performance was similar for the stochastic C-VAE and Diffusion models, the Diffusion policy generalized better, successfully completing $3.6\%$ more unseen test sequences than the other two. This highlights the advantage of its capacity to model complex, multi-modal action distributions for robust generalization. In addition, we observe a dramatic performance drop when training on offline data, highlighting the importance of self-play for learning to recover from mistakes (Ross et al., 2011).
Long-Horizon Sparse Object Goal Chaining: We further test the policies on a challenging long-horizon task (Table 3) where MaskedManipulator receives a sparse object pose goal 2 seconds into the future. Upon reaching the target time, a new 2-second future object goal is set, repeating until the reference motion ends. This evaluates the policy’s ability to reason about distant object states while producing appropriate full-body motion and stable grasps for diverse objects. Performance is measured by the success rate in achieving the sequence of object goals. Consistent with the teleoperation task and prior work (Tessler et al., 2024a), modeling solution diversity proved crucial. While the C-VAE performed better on the training set (indicating some overfitting), the Diffusion model again exhibited superior generalization, successfully solving $9.3\%$ more unseen test sequences than the C-VAE.
Table 2. ‘Teleoperation’: We compare the various architectures on a teleoperation task. Here, MaskedManipulator is conditioned on the positions and rotations of the head, wrists, and object.
Train Test
Success MPJPE [mm] Success MPJPE [mm]
Deterministic 73.8% 16.4 54.6% 24
C-VAE 78.5% 12.7 54.6% 24.4
Offline Diffusion 50% 27.8 25.5% 38.5
[…]
Table 3. Object goals: The agent is provided a sparse objective indicating where (and when) it should transport the object to.
Train Test
Success Tracking Length [s] Success Tracking Length [s]
Deterministic 54.9% 7.8 46.1% 6.2
C-VAE 64.7% 8.3 50.3% 6.6
Offline Diffusion 30.9% 6.8 32.6% 5.8
[…]
Qualitative analysis.
We qualitatively showcase MaskedManipulator’s versatility and the nuanced behaviors it can generate. Figure 6: When provided with goals for wrist position and rotation, MaskedManipulator demonstrates stable and accurate pose tracking. Notably, even without explicit instructions to do so, the policy often infers the appropriate moments to make or break contact with an object to facilitate the target poses, highlighting its learned understanding of physical interaction. Figure 7: MaskedManipulator successfully achieves sparse object goals specified several seconds into the future. For example, it can be tasked to transfer an object to a target location, and the policy will generate the necessary full-body motion, including transferring the object between hands, to complete the objective. Figure 8: When deployed with minimal or no explicit goal constraints, MaskedManipulator exhibits interesting emergent behaviors that reflect the underlying distribution of the GRAB training data. An example of this is the policy picking up a toy airplane and “flying” it through the air, a common interaction pattern present in the dataset for that specific object. This demonstrates its capacity to generate contextually relevant, albeit unprompted, human-like interactions.
These qualitative examples underscore MaskedManipulator’s ability to not only follow specified goals but also to generate rich, physically grounded, and contextually appropriate behaviors learned from the human demonstration data.
7. Limitations
While our framework demonstrates significant progress in versatile loco-manipulation, we identify several limitations that offer opportunities for future research:
Fidelity of the Unified Tracker (MimicManipulator):
We observed that training MimicManipulator to track the full diversity of the GRAB dataset, while successful, sometimes results in slightly lower reconstruction fidelity for individual complex sequences compared to a tracker specialized (i.e., overfit) to only that single motion. This suggests that the current policy representation or capacity might be a bottleneck when learning a very broad repertoire of intricate skills. Future work could explore more expressive policy architectures or adaptive mechanisms within the tracker.
Reconstruction Coverage:
Despite our efforts in data processing and robust reward design, MimicManipulator does not reconstruct all motions from the processed GRAB dataset. Certain interactions remain challenging for the current system to reproduce faithfully within the physics simulation, and the performance of MaskedManipulator on new and unseen tasks leaves room for improvement.
Granularity of Control in MaskedManipulator:
Currently, MaskedManipulator infers a significant amount of the interaction behavior (e.g., specific grasp points on an object, precise timing of contact engagement) based on the sparse goals and its learned priors from MimicManipulator. It does not allow for explicit user control over how an interaction should be performed at a fine-grained level, such as specifying exact contact locations on the object surface.
Addressing these limitations could lead to more capable, precise, and controllable virtual humans, further bridging the gap between high-level creative intent and realistic, interactive animation.
8. Conclusions
In this work, we presented a novel two-stage framework for versatile, physics-based loco-manipulation, enabling digital humanoids to perform complex interactions with objects in response to intuitive, high-level goals. Our approach bridges the gap between the need for precise physical execution and the desire for flexible, versatile control. The resulting MaskedManipulator successfully synthesizes diverse and physically plausible full-body behaviors from sparse, underspecified goals, effectively navigating the multi-modal solution spaces inherent in versatile control.
Our experiments demonstrate that this framework can produce sophisticated manipulation behaviors, such as precise object placements, hand-to-hand transfers, and dynamic interaction with tools, all driven by simple goal specifications. The ability to control both the character and the object through a unified interface opens up exciting possibilities for animators and creators of interactive virtual experiences, simplifying the creation of nuanced and believable character actions.